THE RESURRECTION OF THE HIT EMP DATABASE AND CURRENT UPDATES FOR THE HIT RAN DATABASE

\underline{I. E. GORDON}, L. S. ROTHMAN, Harvard-Smithsonian Center for Astrophysics, Atomic and Molecular Physics Division, Cambridge MA 02138-1516, USA.

Unlike the HIT RAN database, the analogous companion database that is designed for use at high temperatures (HIT EMP) has not evolved during the last decade and therefore was not up-to-date with respect to recent measurements, theoretical calculations, and consistency with HIT RAN^{a}. HIT EMP is currently being resurrected, and the spectral parameters of the H_{2}O, CO_{2}, CO, NO, OH, and CH_{4} molecules are being updated and included in the database. Clearly experimental and theoretical treatments are very challenging in order to achieve accurate line parameters covering wide spectral regions and a great number of vibration-rotation bands. Considerable care is being taken to find the best available values for HIT EMP. The process of revitalizing the HIT EMP database will be presented. In addition, some current updates to the HIT RAN database will be presented, and further improvements towards HIT RAN2008 will be discussed.

aThe update of HIT EMP database has been supported by the CHEMS (Computation of Highly Excited Molecular Spectra)

SBIR project through Spectral Sciences, Inc. HIT RAN database is supported by the NASA Earth Observing System (EOS) under the grant NAG5-13524.